# Evaluating Scoring Bias in LLM-as-a-Judge

### v3

@@ Introduction
Secondly, mitigating scoring bias needs further exploration. Although there are no dedicated methods for addressing scoring bias, intuitive and straightforward approaches, such as scoring multiple times followed by majority voting or averaging the scores, require empirical validation for their effectiveness in mitigating scoring bias. Moreover, the elaborate design of scoring prompt templates, the construction of debiased scoring-based judgment datasets (Park et al. 2024), and adversarial training are potential approaches to enhancing the scoring robustness of judge models.
Additionally, although our work demonstrates the existence of scoring bias in LLM-as-a-Judge, the underlying causes of scoring bias remain to be validated. Approaches such as training data analysis and information flow observation (Yuan et al. 2025) may help identify the reasons for scoring bias, whether it originates from within the model or from external factors.
Conclusion
This work investigates scoring bias in LLM-as-a-Judge, where scoring judgments shift due to minor perturbations in judge prompts. To systematically evaluate this, we designed a framework that perturbs score rubric order, score IDs, and reference answers, quantifying the bias using scoring accuracy and scoring tendency. Experiments on synthesized datasets expanded from four benchmarks confirm that scoring bias significantly undermines the robustness and reliability of existing LLM judges, with varying sensitivity across different models. To mitigate this, we recommend unconventional prompt designs—such as descending rubric order or using letter or Roman numeral IDs—and providing a full-mark reference answer when applicable, both of which effectively improve performance. In summary, our work provides systematic evidence and actionable recommendations for evaluating and enhancing the reliability of LLM judges.

### v4

@@ 5 Limitations and Future Work
Secondly, mitigating scoring bias needs further exploration. Although there are no dedicated methods for addressing scoring bias, intuitive and straightforward approaches, such as scoring multiple times followed by majority voting or averaging the scores, require empirical validation for their effectiveness in mitigating scoring bias. Moreover, the elaborate design of scoring prompt templates, the construction of debiased scoring-based judgment datasets [29], and adversarial training are potential approaches to enhancing the scoring robustness of judge models.
Additionally, although our work demonstrates the existence of scoring bias in LLM-as-a-Judge, the underlying causes of scoring bias remain to be validated. Approaches such as training data analysis and information flow observation [49] may help identify the reasons for scoring bias, whether it originates from within the model or from external factors.
6 Conclusion
This work investigates scoring bias in LLM-as-a-Judge, where scoring judgments shift due to minor perturbations in judge prompts. To evaluate this, we designed a framework that perturbs score rubric order, score IDs, and reference answers, quantifying the bias using scoring stability, scoring accuracy and scoring tendency. Experiments on synthesized datasets expanded from four benchmarks confirm that scoring bias significantly undermines the robustness and reliability of existing LLM judges, with varying sensitivity across different models. To mitigate this, we recommend unconventional prompt designs—such as descending rubric order or using letter or Roman numeral IDs—and providing a full-mark reference answer when applicable, both of which effectively improve performance. In summary, our work provides evidence and actionable recommendations for evaluating and enhancing the reliability of LLM judges.
{credits}
[…]
The authors have no competing interests to declare that are relevant to the content of this article.